OpenStudy (anonymous):

Ann sells bracelets for $4 each and necklaces for $8 each. Which inequality shows x, the number of bracelets and y, the number of necklaces Ann must sell to make at least $100? 4x+8y<= 100 4x+85>= 100 8y + 4x <= 100 8y+4x>=100 But arent b and d basically the same thing? wouldnt either word and mean the same exact thing

OpenStudy (anonymous):

for b you wrote \[ 4x+85\geq 100\] and for d \[8y+4x\geq100\] and they are not the same

OpenStudy (anonymous):

oops !! i meant 4x + 8y \[4x + 8y \ge 100\]

OpenStudy (anonymous):

if you meant \[4x+8y\geq100\] and \[8y+4x\geq 100\] then yes, they are the same

OpenStudy (anonymous):

So either would be right?